# Planning for Robust Open-loop Pushing: Exploiting Quasi-static Belief Dynamics and Contact-informed Optimization

### v1

@@ 8 Discussion
Furthermore, we have shown that informed prior distributions for sampling candidate actions is beneficial, if not necessary, for sampling-based optimization for contact-rich manipulation. We used the product of Gaussian priors to bias the sampling towards smooth and contact-making trajectories. As the evolutionary optimization algorithm CMA-ES is based on sampling from and iterating on Gaussian distributions, we incorporated our Gaussian-distributed prior to initialize CMA-ES. Yet, we see great potential in the use of non-Gaussian priors that are further optimized with a stochastic optimization algorithm such as BS-VP-STO. Especially when the system has many degrees of freedom such as for two robot arms or articulated robotic hands, sampling from a proposal distribution that captures possibly non-Gaussian correlations between the degrees of freedom, e.g. correlations between fingers, is expected to be a key to scalable, realtime control through contacts.
Last, we show that a manipulation task such as pushing, which is typically approached with high-bandwidth closed-loop control, can also be stabilized by planning appropriate open-loop actions that deliberately optimize for robustness. However, if the robot is not able to anticipate perturbations or the statistical properties of the perturbations, feedback is required to be able to stabilize the manipulation. This opens the question of the optimal combination of open-loop robust planning and continuous feedback. Closing the loop may be done by re-planning robust actions if the observations deviate from the planned belief, e.g. in a low-bandwidth feedback loop.

### v2

@@ 8 Discussion
Furthermore, we have shown that informed prior distributions for sampling candidate actions is beneficial, if not necessary, for sampling-based optimization for contact-rich manipulation. We used the product of Gaussian priors to bias the sampling towards smooth and contact-making trajectories. As the evolutionary optimization algorithm CMA-ES is based on sampling from and iterating on Gaussian distributions, we incorporated our Gaussian-distributed prior to initialize CMA-ES. Yet, we see great potential in the use of non-Gaussian priors that are further optimized with a stochastic optimization algorithm such as BS-VP-STO. Especially when the system has many degrees of freedom such as for two robot arms or articulated robotic hands, sampling from a proposal distribution that captures possibly non-Gaussian correlations between the degrees of freedom, e.g. correlations between fingers, is expected to be a key to scalable, real-time control through contacts.
Last, we show that a manipulation task such as pushing, which is typically approached with high-bandwidth closed-loop control, can also be stabilized by planning appropriate open-loop actions that deliberately optimize for robustness. However, if the robot is not able to anticipate perturbations or the statistical properties of the perturbations, feedback is required to be able to stabilize the manipulation. We thus show that the developed approach can be used in a closed-loop controller by integrating the open-loop planning algorithm in a receding-horizon control scheme and by feeding back observations of the object configuration. We furthermore show that a model-predictive controller based on receding-horizon BS-VP-STO even outperforms a controller based on a deterministic model for pushing tasks.